A and B working together can finish a piece of work in 20 days. After working together for 4 days, A quits and B finishes the work in 24 days more. Find the number of days it requires each of them to do the work alone.
Answer by htmentor(1343)

Their combined rate of work is 1 job/20 days = 1/20 job/day
After working together for 4 days, they have completed (1/20)*4 = 1/5 of the job
B finishes the remaining 4/5 of the job in 24 days, so B's rate of work is 4/5 job/24 days = 1/30 job/day
Let r = A's rate of work. Their combined rate = 1/20 = 1/30 + 1/r
Solve for r:
1/r = 3/60 - 2/60 = 1/60 job/day
So B can complete the job in 30 days, A can complete the job in 60 days